OpenStudy (anonymous):

Write the partial fraction decomposition of: 4/2x^2-5x-3

OpenStudy (anonymous):

I know the answer is 4/7(x-3) - 8/7(2x+1). Just do not understand how they get the answer.

OpenStudy (anonymous):

Not having an X value in the numerator confuses me

OpenStudy (anonymous):

so i get 4 = A(2X+1) + B(X-3)

OpenStudy (nocipher):

Ack, actually try X=3... that cancels out the B variable and then X=1/2 which will cancel out the A variable... that should make the equations easy.

OpenStudy (nocipher):

or rather X=-1/2 for the second... sorry, lots of typos tonight

OpenStudy (anonymous):

so 0=2A+B or 1/2=2A+B ?

OpenStudy (anonymous):

-1/2 oops

OpenStudy (nocipher):

If you use X=3 and X=-1/2 you should get 4 = A(2*3+1) + B(3-3) and 4 = A(2(-1/2)+1) + B((-1/2)-3) respectively

OpenStudy (anonymous):

so basically i set x to equal a value that zeros one of the variables out?

OpenStudy (nocipher):

Precisely.

OpenStudy (nocipher):

You can technically pick any two different values of x for this problem. Certain choices make the resulting linear system of equations easier though.

OpenStudy (anonymous):

Ok I will try this. I really appreciate your help.